OpenStudy (anonymous):

Need some help solving this radical equation: ^2√(x−5) = 2

OpenStudy (anonymous):

\[\sqrt[2]{x-5} = 2^{2}\]

OpenStudy (anonymous):

\[\sqrt[2]{x-5} = 2^{2}\]

OpenStudy (anonymous):

@elite ?

OpenStudy (anonymous):

@ParthKohli ?

OpenStudy (anonymous):

@Safa102 @mayankdevnani

OpenStudy (anonymous):

someone please help @mathstudent55 @Hayhayz @imqwerty @Luigi0210 @SolomonZelman

OpenStudy (serenity74):

Solving for X?

OpenStudy (anonymous):

yes! and identifying if it is an extraneous solution

OpenStudy (serenity74):

Square it on both sides

OpenStudy (anonymous):

that's where i got stuck because of the \[\sqrt[2]{x-5}\] i dont know how to square it with the index being 2 like that

OpenStudy (serenity74):

It's still a square root lol so you can square it on both sides you will get x -5 = 2^4 (i am pretty sure it's the same lol)

OpenStudy (anonymous):

oh it's the same? so x-5 = 4 +5 +5 x = 9?

OpenStudy (anonymous):

but my options are: x = 6, solution is not extraneous x = 6, solution is extraneous x = 11, solution is not extraneous x = 11, solution is extraneous

OpenStudy (serenity74):

2 to the power of 4 isn't 8 lol It's 16 x-5 = 16 x = 11

OpenStudy (serenity74):

I am not sure which one it is though. It could be D or C.

OpenStudy (anonymous):

oh sorry i typed the equation wrong it reads \[\sqrt[2]{x-5} = 2\]

OpenStudy (anonymous):

i was looking at the second step i had written down and accidentally added 2^2

OpenStudy (serenity74):

Ohhhh x -5 = 4 x = 9 but that's not one of the options....hmmm

OpenStudy (anonymous):

is the ^2 at the beginning multiplying the equation?

OpenStudy (phi):

can you post a screen shot of the original equation ?

OpenStudy (serenity74):

Well @phi is here you don't need me anymore lol.

OpenStudy (anonymous):

thank you @Serenity74 for your help ! @phi here is the screenshot

OpenStudy (phi):

Lousy font. I would think the first 2 was meant to show square root, but it is the same size and shape as the 2 on the other side. Also, we won't get an answer that matches the choices if we assume the 2 means square root. So they probably mean \[ 2 \sqrt{x-5}= 2\]

OpenStudy (phi):

Though you could first square both sides (to get rid of the square root sign), I would first divide both sides by 2

OpenStudy (anonymous):

Oh that is true! And alright, if I squared both sides would I still get the same answer if I divided?

OpenStudy (phi):

yes, if you don't make a mistake what do you get ?

OpenStudy (anonymous):

\[\sqrt{x-5}\]

OpenStudy (anonymous):

did i delete something? im really bad at using openstudy sorry

OpenStudy (anonymous):

but i think i got just the squareroot x -5 =1

OpenStudy (phi):

if you divide by 2 on both sides you would get \[ \frac{2}{2} \sqrt{x-5}= \frac{2}{2} \] and 2/2 is 1 so that becomes \[ \sqrt{x-5}=1 \] now square both sides

OpenStudy (phi):

I assume you know 1*sqrt(x-5) is just sqrt(x-5) because 1 times anything is the anything

OpenStudy (phi):

once you get \[ \sqrt{x-5}= 1\] squaring the left side gets rid of the radical on the right 1*1 is 1

OpenStudy (anonymous):

yes! I finished the rest of the equation and got x=6. Can I ask, why does nothing happen to the radical when you divide 2 on both sides?

OpenStudy (phi):

think of the radical (for the moment) as just some number (pretend it's 3) then you are asking what happens when we divide \[ 2 \cdot 3 \] by 2 dividing by 2 is the same as multiplying by ½ so it would be \[ \frac{1}{2} \cdot 2 \cdot 3 \] when you multiply you can do it in any order. we can do the ½ * 2 first \[ \left( \frac{1}{2}\cdot 2\right) \cdot 3 \] and ½ * 2 is 1 \[ 1 \cdot 3\] or just 3 the same thing happens if 3 were sqrt(x-5)

OpenStudy (phi):

notice that is different from *adding* for example 2+3 divided by 2 is not so nice: \[ \frac{2+3}{2} = \frac{2}{2} + \frac{3}{2} \] and you see we have to divide both terms by 2.

OpenStudy (phi):

back to your problem. check that 6 "works" in the original equation. If it does, it is *not* an extraneous solution. in this case 2*sqrt(6-5) = 2 2*sqrt(1) = 2 2*1 = 2 2=2 it works.

OpenStudy (anonymous):

thank you for explaining, that makes perfect sense now! and i was just about to ask about the extraneous thing. I thought you had to multiply 2 by 6 and 5, then subtract. And does multiplying the square root by 2 cancel out the radical symbol?

OpenStudy (phi):

do you remember order of operations? PEDMAS (or whatever acronym you might have seen) the square root is treated like an exponent (½) so to evaluate \[ 2 \sqrt{6-5} \] think of it as \[ 2 \cdot (6-5)^\frac{1}{2} \] first do the stuff in parens then do the exponents then the multiply ***And does multiplying the square root by 2 cancel out the radical symbol?*** no, only "squaring" cancels out a square root

OpenStudy (phi):

so the first thing you do is 6-5 and get 1 \[ 2 \cdot 1^\frac{1}{2} \] 1 to *any power* is still 1 (so we can do that 1^½ in our head... we get 1) \[ 2 \cdot 1 \] finally 2 times 1 is 2

OpenStudy (anonymous):

oh okay! thank you so much for your help, you're a lifesaver. My school doesn't do the best job at explaining mathematics. I do have two more to go and would really appreciate if you would help :) The first one is just some help on checking to see if the solution is extraneous and the second is a "see where eloise went wrong" type question.

OpenStudy (phi):

please make a new post.

OpenStudy (anonymous):

will do! :)